OpenStudy (mallorysipp234):

Identify the 17th term of a geometric sequence where a1 = 16 and a5 = 150.06. Round the common ratio and 17th term to the nearest hundredth. a17 ≈ 123,802.31 a17 ≈ 30,707.05 a17 ≈ 19,684.01 a17 ≈ 216,654.05

OpenStudy (mallorysipp234):

@satellite73 can you help

OpenStudy (johnweldon1993):

This one was weird....but it did come out to one of your choices lol so to find the common ratio we have 150.06 x 16r^4 r^4 = 150.06 ------ 16 \[\large r = \sqrt[4]{\frac{150.06}{16}}\] Then round that...what do you get?

OpenStudy (mallorysipp234):

I got 1.7499

OpenStudy (johnweldon1993):

And rounded to the nearest hundredth...?

OpenStudy (mallorysipp234):

im so confused

OpenStudy (johnweldon1993):

Well you get 1.7499 The question says to round the ratio to the nearest hundredth....that would be 1.75 right?

OpenStudy (mallorysipp234):

yes, but that's none of the answer choices....

OpenStudy (johnweldon1993):

That's because we are not done yet...

OpenStudy (johnweldon1993):

We need to find the 17th term of this sequence... To do that...we take our first term...(16) and multiply it by the common ratio ...but raised to the 16th power \[\large 16 \times 1.75^{16}\] what do you get then?

OpenStudy (mallorysipp234):

123802

OpenStudy (johnweldon1993):

123,802.31...but yes.. so answer choice A

OpenStudy (mallorysipp234):

@johnweldon1993 can you help me with this one too??

OpenStudy (mallorysipp234):

@johnweldon1993